Height of Australian women is normally distributed with a mean of 172 cm and a standard deviation of 8 cm. The 5% tallest Australian women would have a height at least ___ cm


1
Expert's answer
2021-01-10T17:26:50-0500

"Mean, \mu = 172"

"Standard\ Deviation, \sigma = 8"


The top "5\%" are the individuals in the region "\alpha=0.05" in a right tail probability distribution.


The corresponding "Z-value = 1.645"



"Z= \frac{x-\mu}{\sigma}"

Therefore


"x= Z*\sigma +\mu"

"= 1.645*8+172=185.16"

The tallest "5\%" are at least "185.16\ cm"